OpenStudy (anonymous):

If the perimeter of a rectangle farm ss 80 m and its length is 25 cm , what is its width

OpenStudy (anonymous):

Perimeter = 2L +2W (for a rectangle) => 80 mm = 2(25) + 2W \[\frac{ 80-50 }{ 2} = W\]

OpenStudy (anonymous):

is it 15 m

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

The perimeter of a square vegetable garden is 24-and-a-half meters.What is the length of each side o the garden.

OpenStudy (anonymous):

24.5 divided by 4

OpenStudy (anonymous):

is it 6.1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 6m